OpenStudy (anonymous):

Use scientific notation to simplify the expression. Give the answer in standard notation. 5.8 × 10^2 ---------- 2.9 × 10^−2

OpenStudy (lukecrayonz):

Divide 5.8 and 2.9, then since it is scientific notation with division, you would take the exponents and SUBTRACT. So since it is negative 2, it would be 2+2, and the answer would be 2*10^4

OpenStudy (anonymous):

5.8 / 2.9 = 2 10^2 / 10^-2 = 10^4 ergo: it all equals 2 x 10^4

OpenStudy (lukecrayonz):

If it was multiplication, you would add the exponents. So if it was 5.8*10^2 TIMES 2.9*10^-2, you would add them, thus making the answer 16.82*10 Since 2+-2 is 0.